Suppose that Jane spends his entire income on good X and Y. The marginal

utilities of both good X and Y are independent of the amount consumed of other

good. The price of X is $2 per unit of X and the price of Y is $3 per unit of Y.

A distribution of the total utilities of goods X and Y consumed is given below.

Distribution of Total Utilities for Good X and Y for Jane.

Number

of Units

Consumed

Total Utility for X Total Utility for Y

1 20 24

2 38 45

3 54 63

4 68 78

5 80 87

6 90 90

Provided Mr. Jane has a monthly income of $ 24. How many units of

Good X and Good Y should Mr. Jane buy?


1
Expert's answer
2021-08-17T10:28:22-0400


budget constraint

"24=2x\n+3y"

consumption bundle(4,6)

units of x"=20+38+54+68=180 \space units"

units of y"=24+45+63+78+87+90=387\space units"